OpenStudy (anonymous):

find the exact area of the region enclosed by the graph of y=3-2x^2, the lines x=1 and x=-1, and the x axis

OpenStudy (anonymous):

use integration

OpenStudy (anonymous):

\[\int\limits_{}^{}3-2x ^{2} = 3x-2/3x ^{3}\]

OpenStudy (anonymous):

But then what do i do? Do i input x=1 x=-1?

OpenStudy (anonymous):

Do you know fundamental theorem: F(b) - F(a) = F(1) - F (-1)

OpenStudy (anonymous):